OpenStudy (anonymous):

An anthropologist finds there is so little Carbon-14 remaining in a prehistoric bone that instruments cannot measure it. This means there is less than 0.2% of the amount of Carbon-14 the bones would have contained when the person was alive. How long ago did the person die? (The constant for Carbon-14 is 0.00012.) Answer 13,411 years 22,491 years 45,020 years 51,788 years

jimthompson5910 (jim_thompson5910):

N = I*e^(-kt) N = I*e^(-0.00012t) 0.002I = I*e^(-0.00012t) 0.002 = e^(-0.00012t) ln(0.002) = -0.00012t ln(0.002)/(-0.00012) = t 51788.4008 = t t = 51788.4008 which rounds to 51,788 years

OpenStudy (anonymous):

Thanks ;)

jimthompson5910 (jim_thompson5910):

you're welcome